A 10g sample of water at boiling point (O^ C) is boiled completely. How much energy was added to the sample


1
Expert's answer
2020-09-29T06:24:18-0400

According to the equation:

Q = m × Hv

where Q2 - boiling energy, m - mass of the sample, Hv - heat of vaporization

As heat of vaporization of water equals 2.26 × 106 J/kg:

Q = 10 g × 2.26 × 106 J/kg = 10 g × 2.26 × 103 J/g = 2.26 × 104 J


Answer: 2.26 × 104 J